# LlamaLens: Specialized Multilingual LLM for Analyzing News and Social Media Content

### v1

@@ 1 Introduction
  • We develop and release a specialized LLM, LlamaLens, which covers 5 broader capabilities, associated with 19 tasks and 52 datasets in three languages: Arabic, English, and Hindi.
@@ 2.2 News and Social Media Analysis
For news and social media analysis there has been research effort with a special focus on fact-checking and harmful content detection Lee et al. (2024); Quelle and Bovet (2024), and news reliability classification Ibrahim (2024). Quelle and Bovet (2024) demonstrated that LLM agents can be employed for fact-checking by retrieving relevant evidence and verifying claims. Similarly,  Ibrahim (2024) explored fine-tuned LLMs, such as Llama-3, to automate the classification of reliable versus unreliable news articles, specifically in Spanish. The Enhancing Perception Hsu et al. (2024) and FACT-GPT Choi and Ferrara (2024) frameworks tackle misinformation by refining fake news explanations through a conversational refinement approach. Similarly, VerMouth Russo et al. (2023) automates social media fact-checking, contributing to broader efforts in combating misinformation. Additionally, the expert recommendation framework Zhang et al. (2024b) leverages a multi-layer ranking system with LLMs, balancing reliability, diversity, and comprehensiveness when suggesting experts for news events. Other initiatives include Botlitica Musi et al. (2024), which identifies propagandistic content in political social media posts, and JSDRV Yang et al. (2024), which focuses on stance detection and rumor verification. In the realm of investigative journalism,  Ali (2024) introduced a tool to retrieve and summarize relevant documents, while  Alonso del Barrio et al. (2024) focused on detecting framing in television program content. Addressing political bias,  Trhlik and Stenetorp (2024) explored bias identification using LLMs. Additionally,  Alonso del Barrio et al. (2024) proposed prompt-engineering LLMs to analyze framing in spoken content from television programs. A comprehensive study was conducted by Zeng et al. (2024), highlighting the use of LLMs in social media applications.
@@ 2.3 Specialized LLMs
Ling et al. (2023) highlighted the importance of developing specialized models for several reasons. One key reason is that, much like humans, acquiring domain expertise and capabilities often requires years of training and experience. Therefore, it is important to train LLMs with domain knowledge to serve professional level usage. In this direction, a recent work by Kotitsas et al. (2024) explored fine-tuning LLMs to improve claim detection. Bao et al. (2024) trained an LLM, FLLM, using curated knowledge with a focus on the business and media domains. For training, they utilized articles published by Fortune Media. The OpenFactCheck framework Wang et al. (2024d) tackles the evaluation of factual accuracy in LLM-generated content. This customizable architecture enables the assessment of both LLM factuality and fact-checking systems, promoting standardized evaluations essential for advancing research on the reliability and factual correctness of LLMs.  Wang et al. (2024a) proposed an explainable fake news detection framework that uses LLMs to generate and evaluate justifications from opposing parties. A defense-based inference module then determines veracity, improving detection accuracy and justification quality, as demonstrated on two benchmarks. In contrast to prior work, our research focuses on developing a specialized model for a wide range of tasks and capabilities for news and social media analysis, representing the first effort in this direction with multilingual capabilities.
@@ 3.1 Data Curation
For dataset curation, we selected key capabilities and their associated tasks, as illustrated in Figure 1, and identified publicly available datasets aligned with these tasks. The choice of languages was influenced by their prominence in the Arabian Peninsula region, making them essential for analyzing linguistic and cultural diversity in the context of social and news media in the region. The initial collection consists of 103 datasets, some of which we excluded due to their different versions (e.g., we selected ArSarcasm-v2 Abu Farha et al. (2021) instead of version 1). After initial pre-selection, the resulting collection consists of 52 datasets as detailed in Tables 3, 4 and 5 in Appendix B. The datasets span various sources, including social media posts, news articles, political debates and transcripts. It consists of $\sim$2.7m samples and a total of 222 labels, reflecting the complexity of tasks such as checkworthiness, claim detection, cyberbullying, emotion detection, news categorization, and more.
@@ 3.2 Preprocessing
After collecting the datasets, we observed that while most were pre-divided into training, development, and test sets, 18 datasets lacked these splits. In cases where datasets were not pre-split, we partitioned them into 70% for training, 20% for testing, and 10% for development. For datasets containing only training and test sets, we further divided the training set, allocating 30% for development. To preserve the class distribution across splits, we employed stratified sampling. We applied several other preprocessing steps such as removing duplicates, unifying labels (e.g., check-worthiness to checkworthiness, fixing uppercase to lowercase), and removed entries with less than 3 letters. These preprocessing steps ensured that the datasets were clean, well-structured, and ready for subsequent analysis or model training. After preprocessing, we obtained a total of 1.2m, 1.4m and 0.14m samples for Arabic, English, and Hindi, respectively. The number of labels in the datasets ranges from 2 to 42. The datasets also include both multiclass and multilabel setups. We provide distribution of the datasets, number of labels and their splits in Tables 3, 4 and 5 in Appendix B. The datasets come in different sizes, ranging from the smallest (e.g., CT-24 subjectivity) to the largest (e.g., English news summarization dataset), and with varying label distributions, from skewed (e.g., propaganda) to more balanced (e.g., Arabic CT22 claim detection). In Appendix A, we provide detailed descriptions of the tasks and datasets.
@@ 4.1 Instruction Dataset
Our approach follows the typical and effective pipeline of aligning LLMs with user intentions and tasks by LLM fine-tuning on representative data. Such approach usually involves creating instruction datasets starting from existing NLP datasets Longpre et al. (2023). An instruction sample is a triplet of a natural language instruction describing the task, an optional input, and an output that represents the answer following the instruction Wang et al. (2023).
@@
There are several potential techniques to create the needed natural language instructions, including manual and automatic approaches. As instructions diversity positively affects model performance and generalization Dubois et al. (2024); Pang et al. (2024); Zhang et al. (2024a), we aim to create a diverse instruction dataset. Due to the scale of tasks and datasets of focus in this work, creating such a diverse set manually is time-consuming and can lead to limited instruction styles Wang et al. (2022); Kaur et al. (2024). We opt to automatically generate instructions by prompting two highly-effective closed LLMs, GPT-4o  [^3] https://platform.openai.com/docs/models/gpt-4o and Claude-3.5 Sonnet,  [^4] https://www.anthropic.com/news/claude-3-5-sonnet to generate a diverse set of 10555We chose the number 10 as a compromise between diversity and the number of samples per instruction. Finding an optimal value requires further study. English instructions per LLM, resulting in 20 instructions per dataset. To ensure the models generate instructions fitting our datasets, we provide the models with the datasets metadata, including dataset name, language, task, task definition and labels space. Exact prompt used to generate instructions and examples of generated instructions can be found in Appendix C. It is important to note that we only require English instructions, as a previous study demonstrated that English prompts outperform language-specific native prompts Kmainasi et al. (2024). Finally, for each input dataset of the 52, we create instructions by appending a randomly selected natural instruction, of the generated 20, per example of each training subset. This guarantees versatile instruction styles even for the same input dataset. Our final instruction tuning dataset is the set of the prepared instructions for all datasets.
@@ 4.2 Training
Our experiments are based on the most effective open LLM to-date, Llama 3.1, which shows effective performance even in multilingual setups Dubey et al. (2024). Fine-tuning larger scales of the model (e.g., 70B version) holds a great overhead in terms of time and computational cost. Moreover, these models might be inaccessible to the greater research community. Thus, we opt to focus on the smallest Llama 3.1-8B version. We particularly base our LlamaLens model on the instruct version of Llama 3.1-8B as it is already aligned to several user tasks.
@@ 4.2.1 Training Setups
We instruction-tune Llama 3.1-8B-Instruct [^6] We use the term Llama-instruct to refer to this model in the rest of the paper. following different setups. Given that fine-tuning LLMs typically requires substantial computational resources, making it a time-consuming and resource-intensive process, therefore, to address this challenge, our main LlamaLens model is fine-tuned in full precision (16-bit) following parameter-efficient fine-tuning using Low-Rank Adaptation (LoRA) Hu et al. (2021). In addition to the full precision model, we aimed to train smaller models, achieving two goals: (i) release smaller but effective models to be used in resource-constrained environments, and (ii) efficiently investigate the effects of some parameter settings on model performance, to guide the full model training. Thus, we also fine-tune the original Llama-8B-Instruct model employing QLoRA Dettmers et al. (2024), which involves quantization of the model’s weights and significantly enhances memory optimization, while maintaining acceptable performance.
Task Dataset Metric SOTA Llama Ours $\Delta$ Task Dataset Metric SOTA Llama Ours $\Delta$
@@
Att.worthiness CT22Attentionworthy W-F1 0.412 0.158 0.544 0.132 Checkworthiness CT24_T1 F1_Pos 0.753 0.404 0.877 0.124
Checkworthiness CT24_T1 F1_Pos 0.569 0.404 0.877 0.308 Claim claim-detection Mi-F1 0.545 0.915
Claim CT22Claim Acc 0.703 0.581 0.778 0.075 Cyberbullying Cyberbullying Acc 0.907 0.175 0.847 -0.060
Cyberbullying ArCyc_CB Acc 0.863 0.766 0.753 -0.110 Emotion emotion Ma-F1 0.790 0.353 0.878 0.088
Emotion Emotional-Tone W-F1 0.658 0.358 0.748 0.090 Factuality News_dataset Acc 0.920 0.654 0.946 0.026
Emotion NewsHeadline Acc 1.000 0.406 0.551 -0.449 Factuality Politifact W-F1 0.490 0.121 0.290 -0.200
Factuality Arafacts Mi-F1 0.850 0.210 0.534 -0.316 News Cat. CNN_News_Articles Acc 0.940 0.644 0.915 -0.025
Factuality COVID19Factuality W-F1 0.831 0.492 0.781 -0.050 News Cat. News_Category Ma-F1 0.769 0.970 0.505 -0.264
Harmfulness CT22Harmful F1_Pos 0.557 0.507 0.508 -0.049 News Genre SemEval23T3-ST1 Mi-F1 0.815 0.687 0.241 -0.574
Hate Speech annotated-hatetweets-4 W-F1 0.630 0.257 0.549 -0.081 News Sum. xlsum R-2 0.152 0.074 0.141 -0.010
Hate Speech OSACT4SubtaskB Mi-F1 0.950 0.819 0.802 -0.148 Offensive Offensive_Hateful Mi-F1 0.692 0.805
News Cat. ASND Ma-F1 0.770 0.587 0.938 0.168 Offensive offensive_language Mi-F1 0.994 0.646 0.884 -0.110
News Cat. SANADAkhbarona Acc 0.940 0.784 0.922 -0.018 Offensive & Hate hate-offensive-speech Acc 0.945 0.602 0.924 -0.021
News Cat. SANADAlArabiya Acc 0.974 0.893 0.986 0.012 Propaganda QProp Ma-F1 0.667 0.759 0.851 0.184
News Cat. SANADAlkhaleej Acc 0.986 0.865 0.967 -0.019 Sarcasm News-Headlines Acc 0.897 0.668 0.956 0.059
News Cat. UltimateDataset Ma-F1 0.970 0.376 0.883 -0.087 Sentiment NewsMTSC Ma-F1 0.817 0.628 0.627 -0.190
News Credibility NewsCredibility Acc 0.899 0.455 0.494 -0.405 Subjectivity CT24_T2 Ma-F1 0.744 0.535 0.508 -0.236
News Sum. xlsum R-2 0.137 0.034 0.075 -0.063 Hindi
Offensive ArCyc_OFF Ma-F1 0.878 0.489 0.652 -0.226 Factuality fake-news Mi-F1 0.759 0.713
Offensive OSACT4SubtaskA Ma-F1 0.905 0.782 0.899 -0.006 Hate Speech hate-speech-detection Mi-F1 0.639 0.750 0.994 0.355
Propaganda ArPro Mi-F1 0.767 0.597 0.762 -0.005 Hate Speech Hindi-Hostility W-F1 0.841 0.469 0.720 -0.121
Sarcasm ArSarcasm-v2 F1_Pos 0.584 0.477 0.307 -0.277 NLI NLI_dataset W-F1 0.646 0.633 0.655 0.009
Sentiment ar_reviews_100k F1_Pos 0.343 0.665 News Sum. xlsum R-2 0.136 0.078 0.117 -0.019
Sentiment ArSAS Acc 0.920 0.603 0.795 -0.125 Offensive Offensive Speech Mi-F1 0.723 0.621 0.847 0.124
Stance stance Ma-F1 0.767 0.608 0.936 0.169 Cyberbullying MC-Hinglish1.0 Acc 0.609 0.233 0.587 -0.022
Stance Mawqif-Arabic-Stance Ma-F1 0.789 0.764 0.867 0.079 Sentiment Sentiment Analysis Acc 0.697 0.552 0.669 -0.028
Subjectivity ThatiAR F1_Pos 0.800 0.562 0.207 -0.593
@@
Table 1: LlamaLens performance across all datasets. SOTA: State-of-the-art results. Llama: Llama 3.1-8B-Instruct. R-2: ROUGE-2, Acc: Accuracy, Mi-F1: Micro-averaged F1, Ma-F1: Macro-averaged F1, W-F1: weighted F1, F1_Pos: F1 of the positive class, –: No SOTA scores found. NLI: Natural Language Inference. Att.worthiness: Attentionworthiness. News Sum.: News Summarization. Cat.: Categorization. : Data was not pre-split.

### v2

@@ 1 Introduction
  • We develop and release a specialized LLM, LlamaLens, which covers 5 broader capabilities, associated with 18 tasks and 52 datasets in three languages: Arabic, English, and Hindi.
@@ 2.2 News and Social Media Analysis
For news and social media analysis there has been research effort with a special focus on fact-checking Shaar et al. (2022), disinformation Hasanain et al. (2023) and harmful content detection Lee et al. (2024); Alam et al. (2022), and news reliability classification Ibrahim (2024). Quelle and Bovet (2024) demonstrated that LLM agents can be employed for fact-checking by retrieving relevant evidence and verifying claims. Similarly, Ibrahim (2024) explored fine-tuned LLMs, such as Llama-3, to automate the classification of reliable versus unreliable news articles. The Enhancing Perception Hsu et al. (2024) and FACT-GPT Choi and Ferrara (2024) frameworks tackle misinformation by refining fake news explanations through a conversational refinement approach. Similarly, VerMouth Russo et al. (2023) automates social media fact-checking, contributing to broader efforts in combating misinformation. Additionally, the expert recommendation framework Zhang et al. (2024b) leverages a multi-layer ranking system with LLMs, balancing reliability, diversity, and comprehensiveness when suggesting experts for news events. Other initiatives include Botlitica Musi et al. (2024), which identifies propagandistic content in political social media posts, and JSDRV Yang et al. (2024), which focuses on stance detection and rumor verification. In the realm of investigative journalism,  Ali (2024) introduced a tool to retrieve and summarize relevant documents, while  Alonso del Barrio et al. (2024) focused on detecting framing in television program content. Addressing political bias,  Trhlík and Stenetorp (2024) explored bias identification using LLMs. Additionally,  Alonso del Barrio et al. (2024) proposed prompt-engineering LLMs to analyze framing in spoken content from television programs. A comprehensive study was conducted by Zeng et al. (2024), highlighting the use of LLMs in social media applications.
@@ 2.3 Specialized LLMs
Ling et al. (2023) highlighted the importance of developing specialized models for several reasons. One key reason is that, much like humans, acquiring domain expertise and capabilities often requires years of training and experience. Therefore, it is important to train LLMs with domain knowledge to serve professional level usage. In this direction, a recent work by Kotitsas et al. (2024) explored fine-tuning LLMs to improve claim detection. Bao et al. (2024) trained an LLM, FLLM, using curated knowledge with a focus on the business and media domains. For training, they utilized articles published by Fortune Media. The OpenFactCheck framework Wang et al. (2025) tackles the evaluation of factual accuracy in LLM-generated content. This customizable architecture enables the assessment of both LLM factuality and fact-checking systems, promoting standardized evaluations essential for advancing research on the reliability and factual correctness of LLMs.  Wang et al. (2024a) proposed an explainable fake news detection framework that uses LLMs to generate and evaluate justifications from opposing parties. A defense-based inference module then determines veracity, improving detection accuracy and justification quality, as demonstrated on two benchmarks. In contrast to prior work, our research focuses on developing a specialized model with a wide range of tasks and capabilities for news and social media analysis, representing the first effort in this direction to incorporate instruction-tuning and multilingual capabilities.
@@ 3.1 Data Curation
For dataset curation, we selected key capabilities and their associated tasks, as illustrated in Figure 1, and identified publicly available datasets aligned with these tasks. Our language choices are influenced by the demographic composition of the Gulf Cooperation Council (GCC) countries, where Arabic is predominant, English serves as a common language, and Hindi is widely spoken due to the significant Indian expatriate population. The initial collection consists of 103 datasets, some of which we excluded due to their different versions. [^3] For example, we have selected ArSarcasm-v2 Abu Farha et al. (2021) instead of version 1. After initial pre-selection, the resulting collection consists of 52 datasets as detailed in Tables 4, 5 and 6 in Appendix B. The datasets span various sources, including social media posts, news articles, political debates and transcripts. It consists of $\sim$2.7m samples and a total of 234 labels, reflecting the complexity of tasks such as checkworthiness, claim detection, cyberbullying, emotion detection, news categorization, and more. In Appendix A, we provide detailed descriptions of the tasks and datasets.
@@ 3.2 Preprocessing
After collecting the datasets, we observed that while most were pre-divided into training, development, and test sets, 18 datasets lacked these splits. In cases where datasets were not pre-split, we partitioned them into 70% for training, 20% for testing, and 10% for development. For datasets containing only training and test sets, we further divided the training set, allocating 30% for development. To preserve the class distribution across splits, we employed stratified sampling. We applied several other preprocessing steps such as removing duplicates, unifying labels (e.g., check-worthiness to checkworthiness, fixing uppercase to lowercase), and removed entries with less than 3 letters. These preprocessing steps ensured that the datasets were clean, well-structured, and ready for subsequent analysis or model training. After preprocessing, we obtained a total of 1.2m, 1.4m and 0.14m samples for Arabic, English, and Hindi, respectively. The number of labels in the datasets ranges from 2 to 42. The datasets also include both multiclass and multilabel setups. We provide distribution of the datasets, number of labels and their splits in Tables 4, 5 and 6 in Appendix B. The datasets come in different sizes, ranging from the smallest (e.g., CT-24 subjectivity) to the largest (e.g., English news summarization dataset), and with varying label distributions, from skewed (e.g., propaganda) to more balanced (e.g., Arabic CT-22 claim detection).
@@ 4.1 Instruction Dataset
Our approach follows the typical pipeline of aligning LLMs with user intentions and tasks by LLM fine-tuning on representative data. Such approach usually involves creating instruction datasets starting from existing NLP datasets Longpre et al. (2023). An instruction sample is a triplet of a natural language instruction describing the task, an optional input, and an output that represents the answer following the instruction Wang et al. (2023).
@@
There are several potential techniques to create natural language instructions, including manual and automatic approaches. As instructions diversity positively affects model performance and generalization Dubois et al. (2024); Pang et al. (2024); Zhang et al. (2024a), we aim to create a diverse instruction dataset. Due to the scale of tasks and datasets of focus in this work, creating such a diverse set manually is time-consuming and can lead to limited instruction styles. We opt to automatically generate instructions by prompting two highly-effective closed LLMs, GPT-4o and Claude-3.5 Sonnet, to generate a diverse set of 10 English instructions [^4] We chose the number 10 as a compromise between diversity and the number of samples per instruction. Finding an optimal value requires further study. per LLM, resulting in 20 instructions per dataset. To ensure the models generate instructions fitting our datasets, we provide the models with the datasets metadata, including dataset name, language, task, task definition and labels space. Exact prompt used to generate instructions and examples of generated instructions can be found in Appendix C. While findings in Kmainasi et al. (2024) show that English prompts generally outperform language-specific counterparts, we adopted a human-centric approach by providing additional native-language instructions to assess the performance of native-instructions after fine-tuning. We followed the same approach above to generate native-instructions. Finally, for each input dataset of the 52, we create instructions by appending a randomly selected natural instruction, of the generated 20, per example of each training subset. This guarantees versatile instruction styles even for the same input dataset. Our final instruction tuning dataset is the set of the prepared instructions for all datasets.
@@ 4.2 Training
We base our experiments on Llama 3.1, the most effective open LLM to date, with strong multilingual performance Dubey et al. (2024). Fine-tuning larger scales of the model (e.g., 70B version) holds a great overhead in terms of time and computational cost. Moreover, These models may be inaccessible to many researchers, so we focus on the smaller Llama 3.1-8B version. We base LlamaLens on Llama 3.1-8B-Instruct, as it is already aligned with various user tasks.
@@ 4.2.1 Training Setups
We instruction-tune Llama 3.1-8B-Instruct [^5] We use the term Llama-instruct or Base to refer to this model in the rest of the paper. following different setups. Given that fine-tuning LLMs typically requires substantial computational resources, making it a time-consuming and resource-intensive process, therefore, to address this challenge, our main LlamaLens model is fine-tuned in bf16 following parameter-efficient fine-tuning using Low-Rank Adaptation (LoRA) Hu et al. (2022). In addition to the full precision model, we aimed to train smaller models, achieving two goals: (i) release smaller but effective models to be used in resource-constrained environments, and (ii) efficiently investigate the effects of some parameter settings on model performance, to guide the full model training. Thus, we also fine-tune the original Llama-8B-Instruct model employing QLoRA Dettmers et al. (2024), which involves quantization of the model’s weights and significantly enhances memory optimization, while maintaining acceptable performance.
Task Dataset Metric SOTA Base L-Lens (Nat.) L-Lens (Eng.) $\Delta$ Task Dataset Metric SOTA Base L-Lens (Nat.) L-Lens (Eng.) $\Delta$
@@
Attention. CT22Attentionworthy W-F1 0.412 0.158 0.454 0.425 0.013 Check. CT24_T1 F1_Pos 0.753 0.404 0.942 0.942 0.189
Check. CT24_T1 F1_Pos 0.569 0.610 0.509 0.502 -0.067 Claim claim-detection Mi-F1 0.545 0.889 0.864
Claim CT22Claim Acc 0.703 0.581 0.756 0.734 0.031 Cyberbullying Cyberbullying Acc 0.907 0.175 0.855 0.836 -0.071
Cyberbullying ArCyc_CB Acc 0.863 0.766 0.833 0.870 0.007 Emotion emotion Ma-F1 0.790 0.353 0.808 0.803 0.013
Emotion Emotional-Tone W-F1 0.658 0.358 0.736 0.705 0.047 Factuality News_dataset Acc 0.920 0.654 0.999 0.999 0.080
Emotion NewsHeadline Acc 1.000 0.406 0.458 0.480 -0.520 Factuality Politifact W-F1 0.490 0.121 0.311 0.287 -0.203
Factuality Arafacts Mi-F1 0.850 0.210 0.738 0.771 -0.079 News Cat. CNN_News_Articles Acc 0.940 0.644 0.970 0.970 0.030
Factuality COVID19Factuality W-F1 0.838 0.492 0.840 0.800 -0.031 News Cat. News_Category Ma-F1 0.769 0.970 0.520 0.824 0.055
Harmfulness CT22Harmful F1_Pos 0.557 0.507 0.535 0.523 -0.034 News Genre SemEval23T3-ST1 Mi-F1 0.815 0.687 0.253 0.241 -0.574
Hate Speech annotated-hatetweets-4 W-F1 0.630 0.257 0.517 0.526 -0.104 News Sum. xlsum R-2 0.152 0.074 0.181 0.182 0.030
Hate Speech OSACT4SubtaskB Mi-F1 0.950 0.819 0.955 0.955 0.005 Offensive Offensive_Hateful Mi-F1 0.692 0.813 0.814
News Cat. ASND Ma-F1 0.770 0.587 0.929 0.919 0.149 Offensive offensive_language Mi-F1 0.994 0.646 0.893 0.899 -0.095
News Cat. SANADAkhbarona Acc 0.940 0.784 0.953 0.954 0.014 Offensive & Hate hate-offensive-speech Acc 0.945 0.602 0.935 0.931 -0.014
News Cat. SANADAlArabiya Acc 0.974 0.893 0.985 0.987 0.013 Propaganda QProp Ma-F1 0.667 0.759 0.973 0.963 0.296
News Cat. SANADAlkhaleej Acc 0.968 0.865 0.982 0.984 -0.002 Sarcasm News-Headlines Acc 0.897 0.668 0.947 0.936 0.039
News Cat. UltimateDataset Ma-F1 0.970 0.376 0.880 0.865 -0.105 Sentiment NewsMTSC Ma-F1 0.817 0.628 0.748 0.751 -0.066
News Cred. NewsCredibility Acc 0.899 0.455 0.933 0.935 0.036 Subjectivity CT24_T2 Ma-F1 0.744 0.535 0.628 0.642 -0.102
News Sum. xlsum R-2 0.137 0.034 0.130 0.129 -0.009 Average 0.773 0.528 0.731 0.747 -0.026
Offensive OSACT4SubtaskA Ma-F1 0.905 0.782 0.882 0.896 -0.009 Hindi
Offensive ArCyc_OFF Ma-F1 0.878 0.489 0.879 0.877 -0.001 Factuality fake-news Mi-F1 0.759 0.993 0.994
Propaganda ArPro Mi-F1 0.767 0.597 0.731 0.747 -0.020 Hate Speech hate-speech-detection Mi-F1 0.639 0.750 0.963 0.963 0.324
Sarcasm ArSarcasm-v2 F1_Pos 0.584 0.477 0.542 0.520 -0.064 Hate Speech Hindi-Hostility W-F1 0.841 0.469 0.753 0.753 -0.088
Sentiment ar_reviews_100k F1_Pos 0.681 0.779 0.785 NLI NLI_dataset W-F1 0.646 0.633 0.679 0.568 -0.078
Sentiment ArSAS Acc 0.930 0.603 0.804 0.800 -0.120 News Sum. xlsum R-2 0.136 0.078 0.170 0.171 0.035
Stance stance Ma-F1 0.767 0.608 0.881 0.926 0.159 Offensive Offensive Speech Mi-F1 0.723 0.621 0.865 0.862 0.139
Stance Mawqif-Arabic Ma f1 0.789 0.764 0.826 0.853 0.065 Cyberbullying MC-Hinglish1.0 Acc 0.609 0.233 0.627 0.625 0.016
Subjectivity ThatiAR F1_Pos 0.800 0.562 0.383 0.441 -0.359 Sentiment Sentiment Analysis Acc 0.697 0.552 0.654 0.647 -0.050
Average 0.773 0.540 0.733 0.735 -0.038 Average 0.613 0.477 0.673 0.656 0.043
@@
Table 1: LlamaLens performance across all datasets. SOTA: State-of-the-art results. L-Lens: LlamaLens, Nat.: Native, Eng.: English. Base: Llama 3.1-8B-Instruct. R-2: ROUGE-2, Acc: Accuracy, Mi-F1: Micro-averaged F1, Ma-F1: Macro-averaged F1, W-F1: Weighted F1, F1_Pos: F1 score for the positive class, –: No SOTA scores found. NLI: Natural Language Inference. Attention: Attentionworthiness. News Cred: News Credibility. News Sum.: News Summarization. Cat.: Categorization. Check: Checkworthiness : Data was not pre-split. The $\Delta$ column represents the difference between LlamaLens (Eng.) and the corresponding SOTA value. Underlined values in the $\Delta$ column indicate cases where LlamaLens (Eng.) outperforms the SOTA.